A number is 72 less than its square. Find all such numbers.
Answer by stanbon(48558)

A number is 72 less than its square. Find all such numbers.
-----------------
Equation:
x^2-x = 72
x^2-x-72 = 0
(x-9)(x+8) = 0
x = 9 or x = -8
-----------------------
If x = 9 then x^2=81 and 81-9 = 72
If x = -8 then x^2 = 64 and 64--8 = 72
================================================
Cheers,
Stan H.